OpenStudy (anonymous):

Which of the following is the equation of a line parallel to the line y = 4x + 1, passing through the point (5,1)?

OpenStudy (ajprincess):

For parallel lines the gradients will be equal. So m=4. \[m=\frac{ y-y_1 }{ x-x_1}\] \[4=\frac{y-1}{x-5}\] \[4(x-5)=y-1\]